# To Buy an Electric Vehicle or Not? A Bayesian Analysis of Consumer Intent in the United States

### v1

@@ Abstract
The adoption of electric vehicles (EVs) is considered critical to achieving climate goals, yet it hinges on consumer interest. This study explores how public intent to purchase EVs relates to four unexamined factors: exposure to EV information, perceptions of EVs’ environmental benefits, views on government climate policy, and confidence in future EV infrastructure; while controlling for prior EV ownership, political affiliation, and demographic characteristics (e.g., age, gender, education, and geographic location). We utilize data from three nationally representative opinion polls conducted by the Pew Research Center between 2021 and 2023, and employ Bayesian techniques to estimate the ordinal probit and ordinal quantile models. Results from ordinal probit show that respondents who are well-informed about EVs, perceive them as environmentally beneficial, or are confident in development of charging stations are more likely to express strong interest in buying an EV, with covariate effects--a metric rarely reported in EV research--of 10.2, 15.5, and 19.1 percentage points, respectively. In contrast, those skeptical of government climate initiatives are more likely to express no interest, by more than 10 percentage points. Prior EV ownership exhibits the highest covariate effect (ranging from 19.0 to 23.1 percentage points), and the impact of most demographic variables is consistent with existing studies. The ordinal quantile models demonstrate significant variation in covariate effects across the distribution of EV purchase intent, offering insights beyond the ordinal probit model. This article is the first to use quantile modeling to reveal how covariate effects differ significantly throughout the spectrum of EV purchase intent.
@@
Decarbonization, electric vehicle, ordinal probit, Pew Research, quantile regression, technology adoption.
journal: arXiv.org
@@ 1 Introduction
Transportation is the largest source of greenhouse gas emissions in the United States (US), representing 28% of the total emissions in 2022, with light-duty vehicles responsible for more than half (Environmental Protection Agency, 2024). Electric vehicles (EVs) are widely regarded as a key solution for reducing emissions and meeting climate targets, given their zero tailpipe emissions. To accelerate EV adoption, governments worldwide, including the US, have implemented a range of policies and incentives. In the US, key measures include nationwide greenhouse gas standards to reduce emissions from new vehicles; up to \$7,500 federal tax credits for EV buyers; investments in charging infrastructure; and state-level mandates like California’s Zero-Emission Vehicle (ZEV) program, which requires automakers to sell an increasing share of EVs [^1] For more details, see the EPA’s greenhouse gas regulations, the IRS EV tax credit guidelines, National Electric Vehicle Infrastructure (NEVI) Program and California’s ZEV Program..
@@
Despite these efforts, EV sales in the US remain relatively low, accounting for only about 10% of global new EV sales in 2023, trailing behind China (60%) and Europe (25%) (International Energy Agency, 2024). At the same time, the future trajectory of EV adoption in the US is increasingly uncertain. In 2024, several automakers, including GM and Ford, scaled back ambitious EV targets, citing weak consumer demand (Colias and Otts, 2024). Moreover, successive administrations have taken sharply different approaches to EVs, creating uncertainty around long-term incentives and regulations. The Biden administration has promoted EV adoption through subsidies, infrastructure investments, and emissions regulations as part of its climate agenda, while the Trump administration has signaled a deregulatory shift on EV policies with plans to weaken both incentives as well as regulations. [^2] President Trump’s Executive Order on January 20, 2025, outlined plans to halt \$7.5 billion in charging infrastructure funding, weaken vehicle emissions standards that drive EV production, eliminate EV tax credits, and revoke California’s EV mandate waiver, which would limit state-level regulations. See https://www.whitehouse.gov/presidential-actions/2025/01/unleashing-american-energy/ These shifting political priorities may reflect broader tensions between environmental regulations, industry interests, and fiscal considerations in EV adoption. While EVs help reduce emissions—particularly in regions with cleaner electricity grids—subsidies and incentives, along with declining gas tax revenue, pose budgetary challenges (Banta et al., 2024). The Trump administration has also cited concerns that stringent emissions and fuel economy regulations increase compliance costs for automakers and restrict consumer choice.
@@
Ultimately, the future of the EV market depends on consumer interest, which is likely to be shaped by multiple factors, including cost considerations, convenience, perception of environmental benefits, and awareness of these trade-offs. Although EVs often have higher upfront costs due to expensive battery technology, they can offer long-term savings through lower operational and maintenance expenses, as electric motors require less maintenance. However, the overall cost of owning an EV can vary significantly depending on factors such as local gasoline and electricity prices, climate, purchase incentives, home charging access, and annual mileage (Woody et al., 2024). Beyond cost considerations, access to reliable charging infrastructure remains an important consideration. Limited charging availability in some regions can exacerbate range anxiety and inconvenience, particularly for long-distance travel. Environmental concerns also shape consumer attitudes, but perceptions may vary, given the regional heterogeneity in short term benefits (Holland et al., 2016) and the perception about critical minerals mining (Lakshman, 2024).
@@
Recognizing the importance of public interest in shaping the future of this market, this study examines how public opinion on purchasing EVs is associated with factors such as exposure to EV information, perceptions about the environmental impact of EVs (as compared to gas-powered vehicles) and government actions to mitigate climate change, confidence in EV related infrastructure development, prior ownership, and demographic characteristics such as age, gender, education, political affiliation, and geographical location of the respondents.
Our paper makes three key contributions. First, we explore the role of several novel factors that have been largely overlooked in the existing EV adoption literature. Previous studies have examined how a combination of factors, including range anxiety, performance concerns, charging infrastructure, environmental attitudes, government policies, peer effects, and demographic characteristics, influences EV purchase intent (Carley et al., 2013; Tiwari et al., 2020; Zhao et al., 2022; Mamkhezri, 2025). Other studies examine the impact of specific factors, such as providing information about total cost of ownership (Dumortier et al., 2015), media coverage (Scherrer, 2023), reputation (Buhmann and Criado, 2023), and consumer incentives (Xue et al., 2023; Stekelberg and Vance, 2024). Related research examines factors driving actual EV sales rather than purchase intent, including tax credits and rebates (DeShazo et al., 2017; Jenn et al., 2018; Clinton and Steinberg, 2019; Sheldon et al., 2023), charging infrastructure (Li et al., 2017), and automakers’ responses to environmental policies (Aghion et al., 2016; Gillingham, 2022; Lohawala, 2023). We contribute to this literature by examining how EV purchase intent is influenced by factors such as confidence in future charging infrastructure development, perceptions of government climate efforts, exposure to EV-related information, and perceptions of EVs’ environmental benefits.
Our second contribution is related to usage of large and nationally representative datasets. Existing research studying the trends and determinants of public opinion and purchase intent for EVs has predominantly relied on small-scale surveys. For example, Tiwari et al. (2020) analyzed 1,800 responses from UK residents, Lashari et al. (2021) analyzed 1,500 responses from South Korea, Hu et al. (2023) analyzed 807 responses from China, and Mamkhezri (2025) analyzed 1,500 responses from the US. [^3] While some recent studies (e.g., Pani et al. (2023) and Ruan and Lv (2022)) use large-scale online sentiment analysis to track the evolution of public perceptions of EVs, they do not examine individual-level predictors of perception. Singh et al. (2020) provide a meta-analysis of 211 peer-reviewed research articles published between 2009 and 2019, identifying key factors influencing public opinion of EVs. Notably, most such studies are either based on a single survey for a particular year, not nationally representative, or constrained by sample size; which diminishes their generalizability. This paper overcomes such data related limitations by drawing on three nationally representative opinion polls conducted by the Pew Research Center in the US with a significantly larger sample size than existing previous studies. Specifically, we use data from 2021, 2022, and 2023, with 11,052, 7,173, and 7,201 observations, respectively.
Our third contribution lies in the application of advanced modeling and estimation techniques to study EV purchase intent. Existing research has used various methods to analyze EV purchase intent, such as linear regression to assess the effects of personal beliefs and preferences (Carley et al., 2013; Krause et al., 2013) and positive media coverage (Scherrer, 2023); binary probit and logit to explore motivational factors like reputation in shaping EV preferences (Buhmann and Criado, 2023); mixed logit models to explore attitudes towards various EV policy incentives, employment changes and electricity costs (Mamkhezri, 2025); structural equation modeling to examine the role of adoption barriers such as range anxiety and charging infrastructure reliability (Tiwari et al., 2020), peer effects (Zhao et al., 2022), and policy incentives (Hu et al., 2023); and topic modeling and sentiment analysis to track shifts in online sentiment toward EVs over time (Ruan and Lv, 2022). These papers rely on Classical (Frequentist) methods, which provide point estimates and confidence intervals, but fail to capture uncertainty in model parameters. This can lead to biased estimates of covariate effects, particularly in non-linear models like probit or logit, where parameters do not directly translate to changes in outcome probabilities (Jeliazkov and Vossmeyer, 2018).
@@
In contrast, this paper adopts a Bayesian modeling framework that provides a posterior distribution over the parameters, systematically incorporates uncertainty in parameters (and observations) when computing covariate effects, and enables us to make direct probability statements about the parameters--rather than relying on the notion of “confidence” as in the Classical framework. Our empirical analysis consists of two stages. First, we estimate an ordinal probit model using Markov chain Monte Carlo (MCMC) methods to assess how different factors are associated with consumer interest in EV adoption. We present the covariate effects on the average probabilities of outcomes, an aspect which is rarely explored in the EV literature. Next, we extend the analysis using ordinal quantile models at the 20th, 50th, and 80th percentiles of purchase intent to examine how different factors shape the decisions of consumers who are more inclined to adopt EVs compared to those who are less inclined. Notably, this study is the first to apply ordinal quantile analysis to EV purchase intent. Since no classical estimation methods exist for ordinal quantile models, Bayesian estimation is essential for this analysis. The results reveal considerable heterogeneity in covariate effects across the distribution of EV purchase intent, which would be missed by ordinal probit models.
@@
The remainder of the article is structured as follows. Section 2 introduces the ordinal probit and ordinal quantile models employed in our analysis, along with their MCMC algorithms for estimation. Section 3 outlines the data and provides a preliminary investigation. In Sections 4, we present the results from our models and discuss their implications. Finally, Section 5 offers concluding remarks.
2 The Modeling Framework
2.1 The Ordinal Probit Model
Ordinal data models (Johnson and Albert, 2000), also known as ordered choice models, allows fitting a regression model to an ordinal dependent (outcome or response) variable, typically denoted by $y$, as a function of the covariates. These models are grounded in the random utility framework from economics and explained, amongst others, in Jeliazkov and Rahman (2012) and Batham et al. (2023). In a typical setting, the response variable $y$ has several categories and each category is assigned a score (value or number) which is inherently ordered or ranked. However, these scores only carry ordinal meaning and lack cardinal interpretation. Thus, while the categories are ranked, the differences between them cannot be compared directly. For example, in our current study, responses to the question about purchasing an EV are coded as follows: 1 for ‘not at all likely’, 2 for ‘not too likely’, 3 for ‘somewhat likely’, and 4 for ‘very likely’. Here, a score of 2 as compared to 1 implies more inclination to buy an electric vehicle, but we cannot interpret a score of 2 as twice the inclination compared to a score of 1. Similarly, the difference in inclination between the scores 2 and 1 is not the same as that between 4 and 3.
In our study, the unobserved variable $z_{i}$ represents intent to purchase EV and is associated with the observed discrete response $y_{i}$, that denotes response to the question about purchasing an EV, as follows:
[…]
where $-\infty=\gamma_{0}<\gamma_{1}\cdots<\gamma_{J-1}<\gamma_{J}=\infty$ are the cut-points (or thresholds) and $J$ denotes the number of categories or outcomes of $y$. This relationship between the latent $z$ (intent to purchase an EV) and observed $y$ (response to the question about purchasing an EV) is visually represented in Figure 1. As seen from the figure, $P(y_{i}=1)$ is the area under $f(z_{i})$ to the left of $\gamma_{1}$, $P(y_{i}=2)$ is the area under $f(z_{i})$ between $\gamma_{1}$ and $\gamma_{2}$, and so on. We set $\gamma_{1}=0$ and $var{(\varepsilon)}=1$ to anchor the location and scale of the distribution, respectively. These restrictions are required for parameter identification and further explained in (Jeliazkov et al., 2008) and Jeliazkov and Rahman (2012).
Figure 1: A pictorial representation of the distribution of the latent variable $z$. The four probabilities P$(y_{i}=1)$, P$(y_{i}=2)$, P$(y_{i}=3)$ and P$(y_{i}=4)$ correspond to the responses: ‘not at all likely’, ‘not too likely’, ‘somewhat likely’, and ‘very likely’, regarding the intention to purchase an electric vehicle ($z_{i}$) for individual $i$ with mean $x^{\prime}_{i}\beta$.
Given the data $y$ = $(y_{1},\cdots,y_{n})^{\prime}$ and the covariates, the likelihood function for ordinal probit model can be written as follows:
where $\Phi(\cdot)$ is the cumulative distribution function (cdf) of a standard normal distribution and $I(y_{i}=j)$ is an indicator function, which equals 1 if the condition within parenthesis is true and 0 otherwise. But the ordering constraint $\gamma_{1}=0<\gamma_{2}<...<\gamma_{J-1}$ is difficult to satisfy during the MCMC sampling. So, we apply the transformation of Albert and Chib (2001): $\delta_{j}=\ln(\gamma_{j}-\gamma_{j-1})$, for $12\leq j\leq J-$ 1, and rewrite the likelihood as a function of $\beta$ and $\delta=(\delta_{2},\ldots,\delta_{J-1})^{\prime}$, which allows us to make inferences from $f(y|\beta,\delta)$.
To complete the Bayesian framework, we utilize the Bayes theorem to combine the augmented likelihood $f(y,z|\beta,\delta)$ with the prior distributions: $\beta\sim N\left(\beta_{0},B_{0}\right)$ and $\delta\sim N\left(d_{0},D_{0}\right)$. The resulting augmented posterior distribution can be written as,
where $f(y_{i}|\delta,z_{i})=1\{\gamma_{j-1}<z_{i}\leq\gamma_{j}\}$, the correspondence between $\gamma$ and $\delta$ is determined by the one-to-one mapping, and the cutpoint index $j$ is determined by the observed value of $y_{i}$. Additionally, $f(z|\beta)=f_{N}(z|X\beta,I_{n})$ and the priors $\pi(\beta)$ and $\pi(\delta)$ are pdf’s of normal distribution. Based on the augmented joint posterior (4), the conditional posteriors can be derived and the model can be estimated using MCMC sampling, as outlined in Algorithm 1.
Post estimation, an important construct is to compute the covariate effects (or marginal effects) on the outcome probabilities, since, unlike linear regression, the coefficients do not give the covariate effect of a variable. Let’s assume that the $l$-th covariate, ${x_{i,l}}$, is set to two distinct values, $a$ and $b$, denoted as ${x_{i,l}^{a}}$ and ${x_{i,l}^{b}}$, respectively. We partition the covariate and parameter vectors as follows: $x_{i}^{a}=(x_{i,l}^{a},x_{i,-l})$, $x_{i}^{b}=(x_{i,l}^{b},x_{i,-l})$, and $\beta=(\beta_{p,l},\beta_{p,-l})$, where $-l$ denotes all covariates (and parameters) except for the $l$-th covariate (and parameter). Our goal is to examine the distribution of the difference ${\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})}$ for $1\leq j\leq J$, by marginalizing over ${x_{i,-l}}$ and $(\beta_{p},\delta_{p})$. The marginalization enables us to incorporate the uncertainty associated with data and model parameters, and avoid nontrivial biases in estimating covariate effects (Jeliazkov and Vossmeyer, 2018).
Algorithm 1: MCMC algorithm for estimating ordinal probit model.
  1. 1.
    Sample $\delta,z|y,\beta$ in one block as follows:
    1. (a)
      Sample $\delta|y,\beta$ marginally of $z$ by generating $\delta^{\prime}$ from a random-walk chain $\delta^{\prime}=\delta+s$, where $s\sim N(0_{J-2},\iota^{2}\hat{D})$, $\iota$ is a tuning parameter, $\hat{D}=\left.-\left[\frac{\partial^{2}\ln f(y|\beta,\delta)}{\partial\delta\partial\delta^{T}}\right]^{-1}\right|_{\delta=\hat{\delta}}$, and $\hat{\delta}=\arg\max_{\delta}\ln f(y|\beta,\delta)$. Accept the proposed value $\delta^{\prime}$ with probability
      otherwise repeat the current value $\delta$.
    2. (b)
  2. 2.
    Sample $\beta|z\sim N(\hat{\beta},\hat{B})$ where $\hat{B}^{-1}=({B}_{0}^{-1}+X^{\prime}X)$, and $\hat{\beta}=\hat{B}(B_{0}^{-1}\beta_{0}+X^{\prime}z)$.
To obtain draws from the distribution ${\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})}$, we use the method of composition (Chib and Jeliazkov, 2006). This involves selecting an individual, extracting the corresponding covariate values, drawing $(\beta,\delta)$ from their posterior distributions, and evaluating the difference ${\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta,\delta)-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta,\delta)}$, where $\Pr(y_{i}=j|x_{i}^{q},\beta,\delta)=\Phi(\gamma_{j}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l})-\Phi(\gamma_{j-1}-x_{i,l}^{q}\,\beta_{l}-x^{\prime}_{i,-l}\,\beta_{-l}),$ for $q=b,a$, and $1\leq j\leq J$. The process is repeated for all individuals and MCMC draws. The average covariate effect (ACE) for outcome $j$ is then computed as the mean of pointwise probability differences:
2.2 The Ordinal Quantile Model
The parameter estimates from the ordinal probit model provide information on the expected value of the latent response variable and the average probabilities of outcomes conditional on the covariates. However, researchers are often more interested in examining specific quantiles of the dependent variable, particularly the lower and upper extremes of the distribution, to answer policy-relevant questions. For instance, having confidence that public charging infrastructure will continue to expand—one of our key covariates—may have less effect on individuals with low interest in EVs. However, among those who are moderately interested, increased confidence in future infrastructure may strengthen intent, reassuring them that charging access will not be a barrier and moving them from “somewhat likely” to “very likely” buyers. To capture how the influence of infrastructure confidence and other factors varies across levels of purchase intent, we use ordinal quantile regression. These insights provide suggestive evidence for more targeted interventions; for example, they show that improving infrastructure visibility may be more effective when directed at consumers already considering EVs, rather than those with little or no interest.
[…]
At this stage, three key issues need attention. First, the ordering constraint of cut-points during sampling is addressed using the transformation, $\delta_{p,j}=\ln(\gamma_{p,j}-\gamma_{p,j-1})$ for $12\leq j\leq J-$ 1, where note that $\delta_{p}$ is quantile dependent. Second, we set $\gamma_{p,1}=0$ to anchor the location and variance is fixed for any quantile to meet the scale restriction. Third, the AL distribution is not convenient for MCMC sampling as it does not yield tractable conditional posteriors. So, following Kozumi and Kobayashi (2011), we utilize the normal-exponential mixture formulation: $\varepsilon_{i}=\theta w_{i}+\tau\sqrt{w_{i}}\,u_{i}$, where $\theta=(1-2p)/(p(1-p))$ and $\tau=\sqrt{2/(p(1-p))}$. The variable $w_{i}$ follows an exponential distribution, $w_{i}\sim\mathcal{E}(1)$, and is independent of $u_{i}$, which follows a normal distribution, $u_{i}\sim N(0,1)$. Based on this formulation, $z_{i}|\beta_{p},w_{i}\sim N(x^{\prime}_{i}\beta_{p}+\theta w_{i},\tau^{2}w_{i})$ for $i=1,\ldots,n$; enabling us to leverage the properties of normal distribution for an efficient MCMC algorithm.
By Bayes theorem, we combine the augmented likelihood $f(y,z|\beta_{p},\delta_{p},w)$ with the normal priors: $\beta_{p}\sim N(\beta_{p0},B_{p0})$, $\delta_{p}\sim N(\delta_{p0},D_{p0})$, to arrive at the augmented joint posterior distribution,
The conditional posterior distributions are derived from the augmented joint posterior distribution (i.e. equation 8), and the parameters are sampled as per Algorithm 2. The implementation of this algorithm, available in the bqror package, is explained in Maheshwari and Rahman (2023).
Algorithm 2: MCMC algorithm for estimating ordinal quantile model.
  1. 1.
    Sample $\beta_{p}|z,w$ $\sim$ $N(\tilde{\beta}_{p},\tilde{B}_{p})$, where,
  2. 2.
    Sample $w_{i}|\beta_{p},z_{i}$ $\sim$ $GIG\,(0.5,\tilde{\lambda}_{i},\tilde{\eta})$, for $i=1,\cdots,n$, where,
  3. 3.
    Sample $\delta_{p}|y,\beta_{p}$ marginally of $w$ (latent weight) and $z$ (latent data), by generating $\delta_{p}^{\prime}$ from a random-walk chain $\delta^{\prime}_{p}=\delta_{p}+s$, where $s\sim N(0_{J-2},\iota^{2}\hat{D})$, $\iota$ is a tuning parameter and $\hat{D}$ denotes negative inverse Hessian, obtained by maximizing the log-likelihood with respect to $\delta_{p}$. Given the current value of $\delta_{p}$, accept the proposed draw $\delta^{\prime}_{p}$ with probability,
  4. 4.
After estimating the ordinal quantile models, we compute the covariate effects using the method of composition. Existing studies on quantile regression that have applied this method to estimate the covariate effects include (Rahman and Vossmeyer, 2019) and Bresson et al. (2021). Assuming a setup similar to Section 2.1, our goal is to obtain draws from the distribution $\{\Pr(y_{i}=j|x_{i,l}^{b})-\Pr(y_{i}=j|x_{i,l}^{a})\}$. We randomly select an individual, extract their covariates, draw a value $(\beta_{p},\delta_{p})$ from their posterior distributions, and finally evaluate $\{\Pr(y_{i}=j|x_{i,l}^{b},x_{i,-l},\beta_{p},\delta_{p})-\Pr(y_{i}=j|x_{i,l}^{a},x_{i,-l},\beta_{p},\delta_{p})\}$, where now for the quantile model: $\Pr(y_{i}=j|x_{i}^{q},\beta_{p},\delta_{p})=F_{AL}(\gamma_{p,j}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})-F_{AL}(\gamma_{p,j-1}-x_{i,l}^{q}\,\beta_{p,l}-x^{\prime}_{i,-l}\,\beta_{p,-l})$, for $q=b,a$ and $1\leq j\leq J$. This process is repeated for all individuals and MCMC draws, and the average covariate effect for outcome $j$ at quantile $p$ is calculated as the mean of difference in pointwise probabilities.

### v2

# Do Determinants of EV Purchase Intent vary across the Spectrum? Evidence from Bayesian Analysis of US Survey Data

@@ Abstract
While electric vehicle (EV) adoption has been widely studied, most research focuses on the average effects of predictors on purchase intent, overlooking variation across the distribution of EV purchase intent. This paper makes a threefold contribution by analyzing four unique explanatory variables, leveraging large-scale US survey data from 2021 to 2023, and employing Bayesian ordinal probit and Bayesian ordinal quantile modeling to evaluate the effects of these variables-while controlling for other commonly used covariates-on EV purchase intent, both on average and across its full distribution. By modeling purchase intent as an ordered outcome-from “not at all likely” to “very likely”-we reveal how covariate effects differ across levels of interest. This is the first application of ordinal quantile modeling in the EV adoption literature, uncovering heterogeneity in how potential buyers respond to key factors. For instance, confidence in development of charging infrastructure and belief in environmental benefits are linked not only to higher interest among likely adopters but also to reduced resistance among more skeptical respondents. Notably, we identify a gap between the prevalence and influence of key predictors: although few respondents report strong infrastructure confidence or frequent EV information exposure, both factors are strongly associated with increased intent across the spectrum. These findings suggest clear opportunities for targeted communication and outreach, alongside infrastructure investment, to support widespread EV adoption.
@@
Decarbonization; Electric vehicle; Ordinal probit; Pew Research Center; Quantile regression; Technology adoption.
journal: Transportation Research Part A
@@ 1 Introduction
Transportation is the largest source of greenhouse gas emissions in the United States (US), representing 28% of the total emissions in 2022, with light-duty vehicles responsible for more than half (Environmental Protection Agency, 2024). Electric vehicles (EVs), given their zero tailpipe emissions, are widely regarded as a key solution for reducing emissions and meeting climate targets (Bhat et al., 2025). To accelerate EV adoption, governments worldwide, including the US, have implemented a range of policies and incentives (Jenn et al., 2018; Mekky and Collins, 2024). Key US measures have included nationwide greenhouse gas standards to reduce emissions from new vehicles; up to \$7,500 federal tax credits for EV buyers; investments in charging infrastructure; and state-level mandates like California’s Zero-Emission Vehicle (ZEV) program, which requires automakers to sell an increasing share of EVs [^1] For more details, see the EPA’s greenhouse gas regulations, the IRS EV tax credit guidelines, National Electric Vehicle Infrastructure (NEVI) Program and California’s ZEV Program. As of late 2025, several of these EV-focused policies have been rolled back under the Trump administration..
@@
Despite these efforts, US sales remain relatively low, accounting for only 10% of global new EV sales in 2023, trailing behind China (60%) and Europe (25%) (International Energy Agency, 2024). The trajectory of US adoption is also increasingly uncertain. In 2024, several automakers, including GM and Ford, scaled back ambitious EV targets, citing weak consumer demand (Colias and Otts, 2024). Successive administrations have taken sharply different approaches to EVs, creating uncertainty around long-term incentives and regulations. The Biden administration promoted EV adoption through subsidies, infrastructure investments, and emissions regulations as part of its broader climate agenda. In contrast, the Trump administration has pursued a deregulatory shift: as of late 2025, the federal EV tax credits are being phased out and the EPA’s endangerment finding is under reconsideration, and California’s Clean Air Act waiver for state-level ZEV mandates has been revoked. [^2] For details on the EV tax credit phase-out, see Section 70502 of the One Big Beautiful Bill Act. For details on the status of California’s waiver for ZEV mandates, see H.J. Res. 88. For details on the reconsideration of the endangerment finding, see the EPA website. Federal charging-infrastructure funding was also suspended but is slated for restoration following litigation (see Department of Transportation newsroom). Other policy changes are subject to ongoing or possible future legal challenges, so further changes remain possible. These changing political priorities reflect tensions between environmental regulations, industry interests, and fiscal considerations in promoting EV adoption. Although EVs reduce emissions—particularly in regions with cleaner electricity grids—subsidies and incentives, along with declining gas tax revenue, pose budgetary challenges (Banta et al., 2024). The Trump administration has also cited concerns that stringent emissions and fuel economy regulations increase compliance costs for automakers and limit consumer choice.
@@
Ultimately, the adoption of EVs depends on consumer interest, which is likely to be shaped by multiple factors, such as cost, convenience, perception of environmental benefits, and awareness of these trade-offs (Rezvani et al., 2015). Although EVs often have higher upfront costs due to expensive batteries, they can offer long-term savings as electric motors require less maintenance (Liu et al., 2021). However, the overall cost of ownership can vary significantly depending on factors such as local gasoline and electricity prices, climate, purchase incentives, home charging access, and annual mileage (Woody et al., 2024). Access to reliable charging infrastructure also remains an important consideration (Chakraborty et al., 2019, 2022). Limited availability in some regions can exacerbate range anxiety and inconvenience, particularly for long-distance travel. Environmental concerns also affect consumer attitudes, but perceptions may vary, given the regional differences in short term benefits (Holland et al., 2016) and the perception about critical minerals mining (Lakshman, 2024).
@@
Recognizing the importance of public interest in shaping the future of this market, we examine how stated purchase intent in the US relates to the following four key factors: exposure to EV information, perceptions of their environmental benefits (compared to gas-powered vehicles), perceptions of government action to mitigate climate change, and confidence in infrastructure development. At the same time, we control for prior ownership, and commonly used demographic variables such as age, gender, education, political affiliation, and geographic location. The data for these four focal factors, along with the dependent and control variables, are drawn from three nationally representative opinion surveys conducted by the Pew Research Center over multiple years.
@@
Our empirical analysis consists of two stages. First, we estimate an ordinal probit model using Markov chain Monte Carlo (MCMC) methods to assess how these factors relate to an ordinal measure of EV purchase intent. We present the covariate effects on the average probabilities of reporting each level of interest, an aspect rarely explored in the EV literature. Secondly, and perhaps more importantly, we extend the analysis using ordinal quantile models at the 20th, 50th, and 80th percentiles of purchase intent to examine how these factors shape the decisions of consumers who are more inclined to adopt EVs compared to those who are less inclined. The results reveal considerable heterogeneity in covariate effects across the distribution of EV purchase intent, which would be missed by ordinal probit models.
Our paper makes three key contributions. First, we examine four unique factors (while controlling for a host of demographic variables) to understand how informational, perceptual, and trust-related factors influence the likelihood and intensity of stated EV purchase intent in the US, and how do these effects vary across different levels of consumer interest. Prior studies of EV purchase intent have considered related factors, but these have either been distinct from the ones we study or investigated only in small samples. The factors we examine have not been analyzed jointly, on large samples, or in the US context. The first factor is confidence in future charging infrastructure development. While earlier work has emphasized the role of existing infrastructure (Carley et al., 2013; Tiwari et al., 2020), purchase intent is also likely to depend on expectations about future network expansion, particularly when the current system remains incomplete. The second factor is perceptions of government climate efforts. Past studies have examined links between purchase intent and political affiliation (e.g., Mamkhezri (2025); Sintov et al. (2020)), but not views on government climate action. Because EVs are widely recognized as a climate policy instrument, individuals who believe the government is already doing enough may be less inclined to purchase—especially if they are less aware of other benefits.
The third factor is exposure to EV-related information. Previous research has examined the effects of information on total cost of ownership (Dumortier et al., 2015), media coverage (Scherrer, 2023), and peer effects (Zhao et al., 2022). We build on this by using a nationally representative survey that directly asks respondents how much they have heard or read about EVs; although subjective, this measure reduces problems of indirect inference. The fourth factor is perceptions of EVs’ environmental benefits. Several studies have considered consumers’ pro-environmental attitudes (Hu et al., 2023; Carley et al., 2013; Krause et al., 2013; Khatua et al., 2023; Scherrer, 2023; Mamkhezri, 2025) and concern about reputation (Buhmann and Criado, 2023). However, research on perceptions of EV’s environmental benefits relative to gasoline powered vehicles is limited. Lashari et al. (2021) and Jia et al. (2025) examine this factor in a small sample, whereas our analysis situates it in a broader US context. [^3] Studies have also examined other factors such as range anxiety, performance concerns, and consumer incentives (See Carley et al. (2013); Tiwari et al. (2020); Zhao et al. (2022); Stekelberg and Vance (2024); Mamkhezri (2025) among others). A separate strand of research examines determinants of actual EV sales, including tax credits and rebates (DeShazo et al., 2017; Jenn et al., 2018; Clinton and Steinberg, 2019; Sheldon et al., 2023), charging infrastructure (Li et al., 2017), peer effects (Chakraborty et al., 2022), and automakers’ responses to environmental policies (Aghion et al., 2016; Gillingham, 2022; Lohawala, 2023). Our findings reveal sizable gaps between the importance of these factors and their prevalence: both infrastructure confidence and information exposure remain low despite their strong association with purchase intent, highlighting opportunities for targeted outreach and investment.
Our second contribution is the use of large, nationally representative datasets, which stands in contrast to the predominantly small-scale, non-representative surveys commonly used in prior research on EV purchase intent. For example, Tiwari et al. (2020), Hu et al. (2023), and Mamkhezri (2025) analyzed 1,800 responses from UK residents, 807 responses from China, and 1,500 responses from the United States, respectively. [^4] Although some recent studies (e.g., Pani et al. (2023) and Ruan and Lv (2022)) use large-scale online sentiment analysis to track the evolution of public perceptions of EVs, they do not examine individual-level predictors of it. Singh et al. (2020) provide a meta-analysis of 211 peer-reviewed research articles 2009–2019, identifying key factors influencing public opinion of EVs. Notably, most such studies are based on single-year, non-representative, or small samples. We overcome these limitations by using three nationally representative US opinion polls, with significantly larger samples: 11,052 in 2021; 7,173 in 2022; and 7,201 in 2023. This allows for greater statistical power and more robust generalizability of our findings.
Our third-and perhaps most significant-contribution lies in applying advanced modeling and estimation techniques to examine the full spectrum of EV purchase intent, addressing the previously unexplored question of whether the determinants of EV adoption vary across different levels of intent. Research has used various methods, such as linear regression to assess the effects of personal beliefs, preferences (Carley et al., 2013; Krause et al., 2013) and positive media coverage (Scherrer, 2023); binary probit and logit to explore motivational factors like reputation (Buhmann and Criado, 2023); mixed logit models to explore attitudes toward various EV policy incentives, employment changes, and electricity costs (Mamkhezri, 2025); structural equation modeling to examine the role of adoption barriers, such as range anxiety and charging infrastructure reliability (Tiwari et al., 2020), peer effects (Zhao et al., 2022), and policy incentives (Hu et al., 2023); and topic modeling and sentiment analysis to track shifts in online sentiment toward EVs over time (Ruan and Lv, 2022). A key limitation of these approaches is that they cannot capture heterogeneity in covariate effects across the full distribution of intent. We address this by applying ordinal quantile analysis to EV purchase intent, an approach not previously used in this context. In contrast to prior work that focuses on average effects, this approach captures heterogeneity in covariate effects both across the full distribution of the intent and across the observed ordinal outcome categories (ranging from “not at all likely” to “very likely” to purchase). This yields more detailed evidence on how different factors shape adoption intent. For example, it shows that confidence in infrastructure development and belief in environmental benefits are not only associated with higher intent among likely adopters, but also with a lower probability of reporting no interest among more skeptical respondents. Note that no classical estimation methods exist for ordinal quantile models, so we use a Bayesian approach.
Finally, our use of a Bayesian framework provides several advantages over traditional (frequentist) methods, which provide point estimates and confidence intervals, but fail to capture uncertainty in model parameters. This can bias covariate effects, particularly in non-linear models like probit or logit, where parameters do not directly reflect changes in outcome probabilities (Jeliazkov and Vossmeyer, 2018). In contrast, Bayesian framework allows us to obtain posterior distribution over the parameters, systematically incorporates uncertainty in parameters (and observations) when computing covariate effects, and enables making direct probability statements about the parameters-rather than relying on the notion of statistical confidence as in the classical framework.
@@
The remainder of the article is structured as follows. Section 2 outlines the data and provides a preliminary investigation. Section 3 introduces the ordinal probit and ordinal quantile models, along with their MCMC algorithms for estimation. In Section 4, we present the results from our models and in Section 5 discuss policy implications. Finally, Section 6 offers concluding remarks.